# Enhancing Chemistry on Quantum Computers with Fermionic Linear Optical Simulation

### v1

@@ IV-B LUCJ-specific Optimizations
for which $e^{-\hat{K}_{2}},e^{\hat{K}_{1}},e^{-\hat{K}_{1}}$ are orbital rotations, $e^{i\hat{J}_{1}}$ is a cluster operator, and $\left|\mathbf{x_{\text{RHF}}}\right\rangle$ is the restricted Hartree-Fock (RHF) state as described by the Jordan-Wigner transformation. The orbital rotation operators $e^{-\hat{K}_{2}},e^{\hat{K}_{1}},e^{-\hat{K}_{1}}$ of $\left|\Psi\right\rangle$ can be fully decomposed into particle number-conserving matchgates. On the other hand, the operator $e^{i\hat{J}_{1}}$ can be fully decomposed into controlled-phase gates.
@@
where $N=\{i|W_{ii}=-1\}$ and $E_{ii}$ is a standard basis matrix. In this sense, any possible $V(x)$ can be written as a base $V_{3}V_{1}$ with a small number of corrections of the form $V_{3}E_{ii}V_{1}$ applied to it. ExtraFerm computes all of these matrices once at runtime. Additionally, ExtraFerm stores the determinants corresponding to each unique pattern of corrections to the base matrix in a hash table to avoid redundant determinant evaluations across trajectories.
@@ V-A Comparisons with Existing Tools
Figure 3: Latency and memory comparisons of different tools when used to calculate the exact probability of an outcome measurement from an arbitrary particle number-conserving extended matchgate circuit. Each circuit contains 16 controlled-phase gates, and the quantity of matchgates grows with circuit size. Results are collected on an Intel Xeon w7-2495X processor. By default, all tools are given access to 24 cores, but only ExtraFerm benefits significantly from using more than 1. We also include results for ExtraFerm restricted to 1 core. Note that the y-axis for each graph is on the logarithmic scale (higher values are orders of magnitude worse).
We emphasize that while LUCJ circuits are an important use case for ExtraFerm, our simulator is a competitive tool for simulating particle number-conserving extended matchgate circuits in general. To demonstrate this, we compared ExtraFerm’s latency and memory consumption against ffsim [ffsim], a state vector simulator that takes advantage of particle number symmetry, and quimb [gray2018quimb], a tensor network simulator. While ffsim computes probabilities for the entire output distribution, quimb is similar to ExtraFerm in the sense that it can compute the probability for a single outcome measurement. We benchmark these three tools on randomly-initialized particle number-conserving extended matchgate circuits ranging from 8 to 36 qubits. These circuits contain 16 controlled-phase gates and a quantity of matchgates proportional to the circuit’s size. Note that these more general circuits are not amenable to the optimizations for LUCJ circuits described in Section IV-B and thus they are not applied.
Fig. 3 demonstrates that for ExtraFerm, latency increases linearly with respect to the number of qubits in the circuit, while for ffsim and quimb, latency increases exponentially. ExtraFerm’s embarrassingly parallel implementation allows it to take full advantage of all 24 CPU cores. On the other hand, ffsim and quimb do not significantly benefit from using more than one core. In terms of memory consumption, system size hardly has an effect on ExtraFerm, while quimb and ffsim suffer an exponential blow-up beginning around the 28-qubit mark.
Figure 4: Speedup as a function of the number of cores on a 24-core Intel Xeon w7-2495X processor for a 40-qubit LUCJ circuit containing 43 controlled-phase gates, with gate angles sampled from $\mathcal{N}(0,10^{-3})$ and total extent 1.732. For each data point, Raw Estimate was used to compute probabilities for 1,000 bitstrings and times were averaged over 10 trials.
Figure 5: Average runtime required to estimate a bitstring probability using Raw Estimate for a 100-qubit LUCJ circuit containing 111 controlled-phase gates, with gate angles drawn from $\mathcal{N}(0,10^{-3})$ and total circuit extent 5.281. Data points were collected for Hamming weights (number of electrons) ranging from 0 to 100 in increments of 2 and averaged over 10 trials. Simulations were performed on a 24-core Intel Xeon w7-2495X processor.
@@ V-B Latency Benchmarks
In this section, we focus on the Raw Estimate algorithm, the fundamental subroutine of the Estimate and an approximate version of the Exact algorithm. Raw Estimate can be run with an arbitrary number of trajectories, making it more general and practically useful. At its core, a Raw Estimate computation amounts to summing contributions from a collection of sampled trajectories. Each trajectory corresponds to a particular decomposition of the controlled-phase gates and contributes an expectation value that must be accumulated into the overall estimate. Accordingly, we examine its runtime behavior, with attention not only to the number of trajectories required but also to non-asymptotic considerations that significantly impact the cost of evaluating each trajectory.
For a given trajectory $x$, the two main costs are constructing the mode transformation matrix $V(x)$ and evaluating the expectation value $\left\langle b\right|V(x)\left|a\right\rangle$. The latter is dominated by determinant calculations, which require selecting an $h\times h$ submatrix of $V(x)$, where $h$ is the Hamming weight of $\left\langle b\right|$ and $\left|a\right\rangle$. In the computational basis, $h$ corresponds to the number of set bits, or equivalently, the number of $X$ gates applied to prepare the Hartree-Fock state in the context of LUCJ circuits. Thus, states with larger Hamming weights—i.e., circuits with more fermions—incur higher computational cost. This can be seen clearly in Fig. 5, in which we fix the angles of a 100-qubit LUCJ circuit and compare the total time required to estimate the probability of a bitstring as a function of the Hamming weight. Independent of the number of trajectories, there is a clear trend indicating higher execution times for circuits for bitstrings with larger Hamming weights. In addition, Fig. 5 shows that for any given Hamming weight, when the number of trajectories is increased by a factor of 10, the respective runtime is also increased by approximately a factor of 10, as we would expect.
However, regardless of the Hamming weight of our bitstring, increasing the number of cores available to the simulator yields near-linear speedups. This is demonstrated in Fig. 4, which shows execution times for running Raw Estimate to estimate the probabilities of 1,000 bitstrings in parallel for different number of trajectories. Furthermore, we can see that as the task becomes harder (more trajectories are used), ExtraFerm scales more effectively.
@@ V-C Error Analysis
Figure 7: Relative error of Raw Estimate versus trajectory count for a 40-qubit LUCJ circuit with 43 controlled-phase gates. These circuits correspond to fermionic systems with 20 spatial orbitals, 2 alpha spin orbitals, and 2 beta spin orbitals. They have extents of 3.484, 62.003, 1,216,201.946, and 8,766,956,281,978.891, respectively (left to right). Orbital rotations are randomly initialized while controlled phase gate angles are sampled i.i.d. from $\mathcal{N}(\theta,10^{-3})$ with $\theta\in\{\pi/64,\pi/16,\pi/4,\pi\}$. Each plot uses the median-probability bitstring from its respective circuit; the bitstring and true probability $p_{\text{true}}$ differ across panels. Blue markers show mean relative error over 10 trials. The horizontal black dashed line marks relative error $\epsilon_{\text{rel}}=.1$; the vertical magenta dashed line is the upper bound on the number of trajectories needed to achieve absolute error $\epsilon_{\text{abs}}=p_{\text{true}}\cdot.1$ with $\delta=10^{-4}$ (using $p_{\max}=p_{\text{true}}$). The green shading highlights the region where both conditions hold: $|\hat{p}-p_{\text{true}}|\leq p_{\text{true}}\cdot.1$ and the trajectory count is at or below the upper bound.
@@
We now share error trends observed when using Raw Estimate to estimate probabilities for LUCJ circuits. We begin with a case study of a 12-qubit LUCJ circuit for which we estimated the probabilities of all outcome measurements in the support with different trajectory counts while varying the average controlled-phase gate angles. In Fig. 6, we display both the observed absolute and relative error for the entire support of the circuit. Across all circuits, we observe that absolute error is proportional to the magnitude of Born-rule probabilities while relative error is inversely proportional to the magnitude of Born-rule probabilities. The fact that a bitstring’s absolute error is proportional to its Born-rule probability ends up being very useful in practice. In section VI, we use this to successfully distinguish high-probability bitstrings from low-probability bitstrings with very small computational overhead.
@@
For all four circuits, we observe decreases in both absolute and relative error as the number of trajectory counts increases. This trend is consistent with the theoretical guarantees of Raw Estimate, which guarantees decreasing additive error as the trajectory count grows. Comparisons between the four circuits are more nuanced. Varying the controlled-phase gate angles alters the Born-rule probabilities of the measurement outcomes. As these angles increase, the output distribution flattens: probability mass spreads across a larger fraction of the support, reducing the dominance of a few high-probability outcomes. This redistribution generally increases both absolute and relative errors. Thus, while increasing trajectory counts consistently reduces error across all circuits, larger controlled-phase angles shift the overall error profile upward.
@@
This phenomenon is illustrated in Fig. 7, which reports the relative error as a function of trajectory count. We examine four 40-qubit LUCJ circuits that again differ only in the angles of their controlled-phase gates. To isolate the effect of these angles on error behavior, we select the median-probability bitstring from the support of each circuit and track its estimated error across increasing trajectory counts. The specific bitstring differs between circuits, but in each case we mark on the $x$-axis the theoretical trajectory count required to achieve a relative error of $0.1$, with the same threshold also indicated by a dashed line on the $y$-axis. For small gate angles ($\theta\approx 0$), this target is reached with relatively few trajectories. As $\theta$ increases, however, the observed errors remain greater than the desired level, and the error curves become steeper, demonstrating a larger difference in performance based on the number of trajectories used.
@@ VI Applications
Figure 8: A schematic of our modification to the SQD algoritm. For the warm-start approach, the green path is taken after the first iteration of configuration recovery, and ExtraFerm is used to calculate Born-rule probabilities of bitstrings that have been corrected to the proper particle number. The gray path is taken for the subsequent iterations of the algorithm. Our experiments demonstrate that this modification improves the accuracy of ground-state energy calculations and reduces the variance between trials.
@@ VI-A Sample-Based Quantum Diagonalization
Sample-Based Quantum Diagonalization (SQD) is an algorithm introduced in Ref. [robledo2025sqd] as an extension of Quantum-Selected Configuration Interaction [kanno2023quantum]. The central idea is to sample chemistry circuits on noisy quantum hardware and then apply classical post-processing to estimate molecular ground-state energies. A key property of LUCJ circuits is that, beyond conserving the total Hamming weight of a bitstring, they also preserve the Hamming weights of the $\alpha$- and $\beta$-spin sectors independently. Equivalently, an LUCJ state can be expressed as $\left|\Psi\right\rangle=\left|\beta\,\alpha\right\rangle$, where the spin sectors must retain their individual Hamming weights after the circuit is applied. SQD corrects bitstrings measured on quantum hardware that violate these conservation rules prior to using them in the energy calculation. Note that SQD is a variational algorithm, which guarantees that the energy estimates produced by it can never be lower than the ground-state energy. In other words, lower energy estimates are better.
We give a brief overview of SQD. First, a molecule is mapped to an LUCJ circuit $\left|\Psi\right\rangle$. Second, this circuit is sampled on a quantum computer to collect $\widetilde{\mathcal{X}}$, a set of noisy bitstrings representing electronic configurations. Third, bitstrings with incorrect particle number are then restored through an iterative procedure called configuration recovery to produce the set of bitstrings $\mathcal{X}_{\text{R}}=\mathcal{X}_{N}\cup\mathcal{X}_{\rightarrow N}$ where $\mathcal{X}_{N}$ and $\mathcal{X}_{\rightarrow N}$ are the bitstrings that originally had the correct particle number and the bitstrings that have been restored to have the correct particle number, respectively. Note that $\mathcal{X}_{\text{R}}$ is not a set in the mathematical sense; duplicate bitstrings are allowed. Fourth, $\mathcal{S}$ is subsampled from $\mathcal{X}_{\text{R}}$ according to observed bitstring frequencies, and the chemistry Hamiltonian is diagonalized in this subspace to obtain an energy value. Lastly, the average orbital occupancies are calculated, which are used to inform the diagonalization in the next iteration of SQD.
@@ VI-B Warm-Start Approach
We propose a simple, cost-effective modification to the SQD workflow that we call warm-starting (see Fig. 8). For the original SQD workflow, $\mathcal{S}$ is subsampled from $\mathcal{X}_{\text{R}}$ according to the empirical frequencies of the bitstrings in $\mathcal{X}_{\text{R}}$. However, due to the large support of these circuits, each corrected bitstring in $\mathcal{X}_{\text{R}}$ appears extremely infrequently for nontrivial chemical systems. This means that $\mathcal{S}$ is subsampled approximately uniformly from $\mathcal{X}_{\text{R}}$, that is, each bitstring is chosen with a probability of $\approx\frac{1}{\#\text{ shots}}$. Since SQD is a variational algorithm, this method of subsampling is beneficial, allowing the solver to explore a diverse set of electronic configurations and converge to the ground state energy.
However, average orbital occupancies identified during early iterations of SQD inform the Hamiltonian diagonalization for later iterations of SQD. Since these bitstrings are chosen randomly, energy estimates can suffer from high variance between trials. Our warm-start approach addresses this by using ExtraFerm to calculate the probabilities of the bitstrings in $\mathcal{X}_{\text{R}}$ during the first iteration of configuration recovery and then choosing the bitstrings with the highest probabilities, rather than subsampling. Later iterations of SQD proceed as usual, subsampling from $\mathcal{X}_{\text{R}}$ according to the empirical frequencies of the corrected bitstrings. We find that the warm-start approach produces energy estimates that are on average more accurate than the ones produced by the original SQD workflow and have lower variance across trials.

### v2

@@ IV-B LUCJ-specific Optimizations
for which $e^{-\hat{K}_{2}},e^{\hat{K}_{1}},e^{-\hat{K}_{1}}$ are orbital rotations, $e^{i\hat{J}_{1}}$ is a cluster operator, and $\left|\mathbf{x_{\text{RHF}}}\right\rangle$ is the restricted Hartree-Fock (RHF) state as described by the Jordan-Wigner transformation. The orbital rotation operators $e^{-\hat{K}_{2}},e^{\hat{K}_{1}},e^{-\hat{K}_{1}}$ of $\left|\Psi\right\rangle$ can be fully decomposed into particle number-conserving matchgates, while the operator $e^{i\hat{J}_{1}}$ can be fully decomposed into controlled-phase gates.
@@
where $N=\{i|W_{ii}=-1\}$ and $E_{ii}$ is a standard basis matrix. In this sense, any possible $V(x)$ can be written as a base matrix $V_{3}V_{1}$ with a small number of corrections of the form $V_{3}E_{ii}V_{1}$ applied to it. ExtraFerm computes all of these matrices once at runtime. Additionally, ExtraFerm stores the determinants of each of these unique patterns in a hash table to avoid redundant determinant evaluations across trajectories.
@@
Summary: Exploiting symmetries in the matchgate approximations of LUCJ circuits allows us to compute probabilities even more efficiently for these circuits.
Figure 4: Latency and memory comparisons of different tools when used to calculate the exact probability of an outcome measurement from a randomly generated particle number-conserving extended matchgate circuit. Each circuit contains 16 controlled-phase gates and a quantity of matchgates proportional to circuit size. Results are collected on an Intel Xeon w7-2495X processor. By default, all tools are given access to 24 cores, but only ExtraFerm benefits significantly from using more than 1. We also include results for ExtraFerm restricted to 1 core. Note that the y-axis for each graph is on the logarithmic scale (higher values are orders of magnitude worse).
@@ V-A Comparisons with Existing Tools
We emphasize that while LUCJ circuits are an important use case for ExtraFerm, our simulator is a competitive tool for simulating particle number-conserving extended matchgate circuits in general. To demonstrate this, we compared ExtraFerm’s latency and memory consumption against quimb [gray2018quimb], a tensor network simulator and ffsim [ffsim], a state vector simulator that takes advantage of particle number symmetry. While ffsim computes probabilities for the entire output distribution, quimb is similar to ExtraFerm in the sense that it can compute the probability of a single outcome measurement. We benchmark these three tools on randomly-initialized particle number-conserving extended matchgate circuits ranging from 8 to 36 qubits. These circuits contain 16 controlled-phase gates and a quantity of matchgates proportional to the circuit’s size. Note that these more general circuits are not amenable to the optimizations for LUCJ circuits described in Section IV-B and thus these optimizations are not applied.
Fig. 4 demonstrates that for ExtraFerm, there is a substantially superior latency scaling over quimb and ffsim. ExtraFerm’s embarrassingly parallel implementation allows it to take full advantage of all 24 CPU cores on the machine that we perform this benchmark on. On the other hand, quimb and ffsim do not significantly benefit from using more than one core. ExtraFerm’s latency scaling advantage for large circuits is still substantial but less extreme when ExtraFerm is restricted to a single CPU core. In terms of memory consumption, system size hardly has an effect on ExtraFerm, while quimb and ffsim suffer an exponential blow-up beginning around the 28-qubit mark.
We note that while ffsim is likely to be state-of-the-art among state vector simulators, our usage of quimb can probably be improved on. For simplicity, we used its exact tensor network contraction functionality, but it might be possible to introduce truncations that leave the result unaffected up to numerical precision. While we placed the controlled-phase gates between random pairs of qubits for our benchmarks, recent work has shown that tensor networks can be effective at simulating certain LUCJ circuits with controlled-phase gates that are geometrically local on a 2-dimensional lattice [rudolph2025simulating].
Figure 5: Speedup as a function of the number of cores used on a 24-core Intel Xeon w7-2495X processor for executing ExtraFerm on a 40-qubit LUCJ circuit containing 43 controlled-phase gates with angles sampled from $\mathcal{N}(0,10^{-3})$ and a total extent 1.732. For each data point, Raw Estimate was used to compute probabilities for 1,000 bitstrings and times were averaged over 10 trials.
@@ V-B Latency Benchmarks
In this section, we focus on the Raw Estimate algorithm, the fundamental subroutine of the Estimate and an approximate version of the Exact algorithm. Raw Estimate can be run with an arbitrary number of trajectories, making it more flexible than Exact. At its core, a Raw Estimate computation amounts to summing contributions from a multiset of sampled trajectories. Each trajectory corresponds to a particular decomposition of the controlled-phase gates and contributes an expectation value that must be accumulated into the overall estimate. Accordingly, we examine its runtime behavior, with attention not only to the number of trajectories required but also to non-asymptotic considerations that significantly impact the cost of evaluating each trajectory.
For a given trajectory $x$, the two main costs are constructing the mode transformation matrix $V(x)$ and evaluating the expectation value $\left\langle b\right|V(x)\left|a\right\rangle$. The latter is dominated by determinant calculations, which require selecting an $h\times h$ submatrix of $V(x)$, where $h$ is the Hamming weight of $\left|a\right\rangle$ and $\left|b\right\rangle$. In the computational basis, $h$ corresponds to the number of set bits, or equivalently, the number of $X$ gates applied to prepare the Hartree-Fock state in the context of LUCJ circuits. Thus, states with larger Hamming weights—i.e., circuits with more fermions—incur higher computational cost. This can be seen clearly in Fig. 6, in which we fix the angles of a 100-qubit LUCJ circuit and compare the total time required to estimate the probability of a bitstring as a function of its Hamming weight. Independent of the number of trajectories, there is a clear trend indicating higher execution times associated with larger Hamming weights. In addition, Fig. 6 shows that for any given Hamming weight, when the number of trajectories is increased by a factor of 10, the respective runtime is also increased by approximately a factor of 10, as we would expect.
However, regardless of the Hamming weight of our bitstring, increasing the number of cores available to ExtraFerm yields near-linear speedups. This is demonstrated in Fig. 5, which shows execution times for running Raw Estimate to estimate the probabilities of 1,000 bitstrings in parallel for different quantities of trajectories. Furthermore, we can see that as the task becomes harder (more trajectories are used), ExtraFerm scales more effectively.
Figure 6: The average runtime required to estimate a bitstring probability using Raw Estimate for a 100-qubit LUCJ circuit containing 111 controlled-phase gates with gate angles drawn from $\mathcal{N}(0,10^{-3})$ and a total circuit extent of 5.281. Data points were collected for Hamming weights (the number of fermions) ranging from 0 to 100 in increments of 2 and averaged over 10 trials. Simulations were performed on a 24-core Intel Xeon w7-2495X processor.
@@ V-C Error Analysis
Figure 8: Relative error of Raw Estimate versus trajectory counts for a 40-qubit LUCJ circuit with 43 controlled-phase gates. These circuits correspond to fermionic systems with 20 spatial orbitals, 2 alpha spin orbitals, and 2 beta spin orbitals. They have extents of 3.48, 62.0, $1.22\cdot 10^{6}$, and $8.77\cdot 10^{12}$, respectively (left to right). Orbital rotations are randomly initialized while controlled phase gate angles are sampled i.i.d. from $\mathcal{N}(\theta,10^{-3})$ with $\theta\in\{\pi/64,\pi/16,\pi/4,\pi\}$. Each plot uses the median-probability bitstring from its respective circuit, so the bitstring and true probability $p_{\text{true}}$ differ across panels. Blue markers show the mean relative error over 10 trials. The horizontal black dashed line marks relative error $\epsilon_{\text{rel}}=.1$ while the vertical magenta dashed line is the upper bound on the number of trajectories needed to achieve absolute error $\epsilon_{\text{abs}}=p_{\text{true}}\cdot.1$ with $\delta=10^{-4}$ (using $p_{\max}=p_{\text{true}}$). The green shading highlights the region where both conditions hold: $|\hat{p}-p_{\text{true}}|\leq p_{\text{true}}\cdot.1$ and the trajectory count is at or below the upper bound.
@@
We now share error trends observed when using Raw Estimate to estimate probabilities for LUCJ circuits. We begin with a case study of a 12-qubit LUCJ circuit for which we estimated the probabilities of all outcome measurements in the support with different trajectory counts while varying the average controlled-phase gate angles. In Fig. 7, we display both the observed absolute and relative error for the entire support of the circuit. Across all circuits, we observe that absolute error is proportional to the magnitudes of Born-rule probabilities while relative error is inversely proportional to the magnitudes of Born-rule probabilities. The fact that a bitstring’s absolute error is proportional to its Born-rule probability ends up being very useful in practice. In section VI, we use this to successfully distinguish high-probability bitstrings from low-probability bitstrings with very small computational overhead.
@@
For all four circuits, we observe decreases in both absolute and relative error as the trajectory count increases. This trend is consistent with the theoretical guarantees of Raw Estimate, which guarantees decreasing additive error as the trajectory count grows. Comparisons between the four circuits are more nuanced. Varying the controlled-phase gate angles alters the Born-rule probabilities of the measurement outcomes. As these angles increase, the output distribution flattens: probability mass spreads across a larger fraction of the support, reducing the dominance of a few high-probability outcomes. This redistribution generally increases both absolute and relative errors. Thus, while increasing trajectory counts consistently reduces error across all circuits, larger controlled-phase angles shift the overall error profile upward.
@@
This phenomenon is illustrated in Fig. 8, which reports the relative error as a function of trajectory count. We examine four 40-qubit LUCJ circuits that again differ only in the angles of their controlled-phase gates. To isolate the effect of these angles on error behavior, we select the median-probability bitstring from the support of each circuit and track its estimated error across increasing trajectory counts. The specific bitstring differs between circuits, but in each case we mark on the $x$-axis the theoretical trajectory count required to achieve a relative error of $0.1$, with the same threshold also indicated by a dashed line on the $y$-axis. For small controlled-phase gate angles ($\theta\approx 0$), this target is reached with relatively few trajectories. As $\theta$ increases, however, the observed errors remain greater than the desired level, and the error curves become steeper, demonstrating a larger difference in performance based on the number of trajectories used.
@@ VI Applications
Figure 9: A schematic of our modification to the SQD algorithm. For the warm-start approach, the green path is taken after the first iteration of configuration recovery, and ExtraFerm is used to calculate the Born-rule probabilities of bitstrings that have been corrected to the proper particle number. The gray path is taken for the subsequent iterations of SQD. Our experiments demonstrate that this modification can improve the accuracy of ground-state energy calculations and reduce variance across trials.
@@ VI-A Sample-Based Quantum Diagonalization
Sample-Based Quantum Diagonalization (SQD) is an algorithm introduced in Ref. [robledo2025sqd] as an extension of Quantum-Selected Configuration Interaction [kanno2023quantum]. The central idea is to sample chemistry circuits on noisy quantum hardware and then apply classical post-processing to estimate molecular ground-state energies. A key property of LUCJ circuits is that, beyond conserving the total Hamming weight of their initial states, they also preserve the Hamming weights of the $\alpha$- and $\beta$-spin sectors independently. Equivalently, an LUCJ state can be expressed as $\left|\Psi\right\rangle=\left|\beta\,\alpha\right\rangle$, where the spin sectors must retain their individual Hamming weights after the circuit is applied. SQD corrects bitstrings measured on quantum hardware that violate these conservation rules prior to using them in the energy calculation. Note that SQD is a variational algorithm, which guarantees that the energy estimates produced by it can never be lower than the ground-state energy. In other words, lower energy estimates are better.
We give a brief overview of SQD. First, a molecule is mapped to an LUCJ circuit $\left|\Psi\right\rangle$. Second, this circuit is sampled on a quantum computer to collect $\widetilde{\mathcal{X}}$, a multiset of noisy bitstrings representing electronic configurations. Third, bitstrings with incorrect particle number are then restored through an iterative procedure called configuration recovery to produce the multiset of bitstrings $\mathcal{X}_{\text{R}}=\mathcal{X}_{N}\cup\mathcal{X}_{\rightarrow N}$ where $\mathcal{X}_{N}$ denotes the bitstrings that originally had the correct particle number and $\mathcal{X}_{\rightarrow N}$ denotes the bitstrings that have been restored to the correct particle number. Fourth, $\mathcal{S}$ is subsampled from $\mathcal{X}_{\text{R}}$ according to observed bitstring frequencies, and the chemistry Hamiltonian is diagonalized in this subspace to obtain an energy value. Lastly, the average orbital occupancies are calculated, which are used to inform the Hamiltonian diagonalization step during the next iteration of SQD.
@@ VI-B Warm-Start Approach
We propose a simple, cost-effective modification to the SQD workflow that we call warm-starting (see Fig. 9). For the original SQD workflow, $\mathcal{S}$ is subsampled from $\mathcal{X}_{\text{R}}$ according to the empirical frequencies of the bitstrings in $\mathcal{X}_{\text{R}}$. However, due to the large support of these circuits, each corrected bitstring in $\mathcal{X}_{\text{R}}$ appears extremely infrequently for nontrivial chemical systems. This means that $\mathcal{S}$ is subsampled approximately uniformly from $\mathcal{X}_{\text{R}}$, that is, each bitstring is chosen with a probability of $\approx\frac{1}{\#\text{ shots}}$. Since SQD is a variational algorithm, this method of subsampling works well, allowing the solver to explore a diverse range of electronic configurations and converge to the ground state energy.
However, because average orbital occupancies identified during early iterations of SQD inform the Hamiltonian diagonalization step during later iterations of SQD, energy estimates can suffer from high variance between trials. Our warm-start approach addresses this by using ExtraFerm to calculate the probabilities of the bitstrings in $\mathcal{X}_{\text{R}}$ during the first iteration of configuration recovery. We then choose the bitstrings with the highest probabilities, rather than subsampling. Later iterations of SQD proceed as usual, subsampling from $\mathcal{X}_{\text{R}}$ according to the empirical frequencies of the corrected bitstrings. We find that the warm-start approach produces energy estimates that are on average more accurate than the ones produced by the original SQD workflow and have lower variance across trials.